OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

help please!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

did I just help you with this stuff? ._. rise over run finds the slope ( y2 - y1 ) / ( x2 - x1 ) ( 1 / 2 ) / ( 2 / -3 ) slope = -.75, or -3/4 slope = m m = -3/4 ( y - y1 ) = m ( x - x1 ) y -1 = -3/4 ( x - (-3) ) balance it y -1 = -3/4 ( x + 3 ) y - 1 = -3/4x - 2.25 y = -3/4x - 1.25 those are really weird numbers, so I'll check my answer after I post this

OpenStudy (anonymous):

@Zilther its either A,B,C,D That isnt the answer

OpenStudy (anonymous):

A.y + 3 = −5(x − 2) B. y − 2 = −5(x + 3) C. y + 3 = −one fifth(x − 2) D. y − 2 = −one fifth(x + 3) @Zilther

OpenStudy (anonymous):

given that the line passes through points ( -3 , 2 ) and ( 2 , 1 ) ( 1 / 2 ) / ( 2 + 3 ) slope = 1/5 y - 3 = 1/5 ( x - 2 )

OpenStudy (anonymous):

I messed up again, wtf y - 2 = 1/5 ( x + 3 )

OpenStudy (anonymous):

the formula is\[( y - y1 ) = m ( x - x1 )\]

OpenStudy (anonymous):

@Zilther LOL IF YOU CANT HELP ME ITS COOL MAN ILL JUST KEEP ASKING

OpenStudy (anonymous):

orz, go ahead, I'll try to check my answers before I post them from now on.